Solve each equation for the variable specified.
(rt + 9)÷5=h , solve for t

Found 2 solutions by Alan3354, etutorworld:
Answer by Alan3354(69443)

(rt + 9)÷5=h , solve for t
(rt + 9)/5=h
rt + 9 = 5h
rt = 5h - 9
t = (5h - 9)/r

Answer by etutorworld(19)

Solve each equation for the variable specified.
(rt + 9)÷5=h , solve for t
To solve for 't' we will have to isolate t from the rest of the variables and constants in the expression
(rt + 9)÷5=h
Multiplying by 5 on both sides
rt + 9 = 5h
Subtracting 9 from both sides
rt + 9 - 9 = (5h -9)
rt = (5h-9)
To get the value in terms of 't' Divide by 'r' on both sides
t = (5h-9)/r [Answer]